OpenStudy (anonymous):

The sum of the digits of a certain two-digit numberis 8. Reversing its digits increases the number by 36. Find the Number.

OpenStudy (dumbcow):

2 digit number has a "tens place" and a "ones place" let x be first digit, y second digit value of number is 10x +y ..... value of reverse order is 10y +x so x+y = 8 and (10x +y) +36 = 10y +x now simplify and solve system of 2 equations

OpenStudy (anonymous):

how do you simplify it?

OpenStudy (dumbcow):

combine like terms so it is in form: Ax +By = C

OpenStudy (anonymous):

ok i got 9x-9y=-36, is that right?

OpenStudy (dumbcow):

correct ... now solve system using elimination (or addition method)

OpenStudy (dumbcow):

wait you could use substitution as well :|

OpenStudy (anonymous):

do we eliminate it using x+y=8?

OpenStudy (dumbcow):

if you are unfamiliar with that method, then just use substitution: solve for y in 1st equation , then plug it into 2nd equation

OpenStudy (anonymous):

ok but im using the equations, x+y=8 and 9x-9y=-36?

OpenStudy (dumbcow):

yes

OpenStudy (anonymous):

is y=12 and x=8?

OpenStudy (dumbcow):

no, x+y = 8 .... 12+8 = 20 ??

OpenStudy (dumbcow):

also a digit is 0-9 .... 12 is 2 digits not one

OpenStudy (anonymous):

y=6 x=2

OpenStudy (dumbcow):

yes, correct so number is 26

OpenStudy (anonymous):

Thank You

OpenStudy (dumbcow):

yw